# Stability of the $L^{p}$-Poincaré inequality for the Lebesgue and Gaussian probability measures with explicit geometric dependence and applications to spectral gaps

### v2

@@ Abstract.
In this paper, we obtain stability results for the $L^{p}$-Poincaré inequality for both Lebesgue and Gaussian probability measures (Theorem 3.3 and Theorem 3.12) that involve explicit dependence on the geometry of the domain. As a byproduct, the explicit constant allows us to recover an important result of Yu, Zhong and Smits (Corollary 3.8), related to the fundamental gap conjecture of the Laplacian (resolved by Andrews and Clutterbuck), thereby providing an alternative proof. Moreover, we extend this spectral gap result to the $p$-Laplacian (Corollary 3.6). Our approach relies on properties of the first eigenfunction of the (Gaussian) $p$-Laplacian operator and weighted Poincaré inequalities for log-concave measures on convex domains.
@@
Poincaré inequality, Stability inequalities, Gaussian Poincaré inequality.
@@ 1. Introduction
and $\langle\cdot,\cdot\rangle$ is the usual scalar product in $\mathbb{R}^{n}$. It is important to note that the result (1.3), in [FT02], was then used to obtain an existence result concerning a (p1)(p-1)-homogeneous problem with Dirichlet $p$-Laplacian.
@@
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx\geq c_{1}(p)\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}d(u,E_{Poin})^{p}.$ (1.5)
@@
$\pi_{p}=2\int_{0}^{+\infty}\frac{1}{1+\frac{1}{p-1}s^{p}}ds=2\pi\frac{(p-1)^{1/p}}{p(\sin(\pi/p))}\quad,\\ \quad c_{1}(p)=\inf_{(s,t)\in\mathbb{R}^{2}\setminus\{(0,0)\}}\frac{\bigl[t^{2}+s^{2}+2s+1\bigr]^{\frac{p}{2}}-1-ps}{\bigl[t^{2}+s^{2}\bigr]^{\frac{p}{2}}}\in(0,1]$
@@
The constant, in (1.5), is explicit with clear dependence on the diameter of the convex domain $\Omega$, which we believe is the first such result. Moreover, after normalizing the eigenfunctions $u_{1}$ and $u_{2}$, we obtain the following spectral gap of the $p$-Laplacian:
$\lambda_{2}(p,\Omega)-\lambda_{1}(p,\Omega)\geq c_{1}(p)\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}C(p,\Omega,u_{1},u_{2}),$ (1.6)
@@
The difference between the first two eigenvalues of the Dirichlet Laplacian on a bounded open domain $\Omega$, denoted by $\lambda_{2}(p,\Omega)-\lambda_{1}(p,\Omega)$, is commonly known as the fundamental gap of $\Omega$. The fundamental gap problem has an interesting history and we give a brief overview of it here. Originally, van den Berg, in [VAN83], first observed that many convex sets have a spectral gap of $\frac{3\pi^{2}}{\text{diam}(\Omega)^{2}}$. This was also independently observed and then conjectured to be true in any convex domain $\Omega$ by Yau [YAU87], Ashbaugh and Benguria [AB89] for the Schrödinger operator $\Delta+V$. The general statement in terms of the Schrödinger operator has been resolved in the one-dimensional setting by Lavine [LAV94] (see, also, [AB89, HOR03]). In higher dimensions, a breakthrough was made by Singer, Wong, Yau and Yau [SWY+85], where they obtained a lower bound of $\frac{\pi^{2}}{4\text{diam}(\Omega)^{2}}$ by utilizing the gradient estimates in spirit of those from Li [LI79], Li and Yau [LY80] and the log-concavity of the first eigenfunction (established for convex potentials by Brascamp and Lieb [BL76]). Later, Yu, Zhong [YZ86] and [SMI96] improved the estimate to $\frac{\pi^{2}}{\text{diam}(\Omega)^{2}}$. Under certain assumptions on the domain $\Omega$, major step forward in proving the conjecture was made by Bañuelos and Méndez-Hernández [BM00], Davis [DAV01] and Bañuelos with Kröger [BK01]. Finally, Andrews and Clutterbuck [AC11] proved the conjecture in 2011 for any convex domain $\Omega$. In a very recent paper [ABF24], Amato, Bucur and Fragalà strengthened the result of Andrews and Clutterbuck and quantified it in terms of flatness, thereby answering an additional open problem by Yau [YAU93]. We also refer to some works related to the fundamental gap problem of the Neumann Laplacian [PW60, ZHO84].
Lastly, we would like to mention that, in general, in most stability results, the explicit form of the constant that appears nearby the distance function (1.1) is unknown. Even for the historic result of Bianchi and Egnell [BE91] for the Sobolev inequality, which was obtained in 1991, no information was known for the constant $c_{S}$ (except an upper bound [KÖN23]) until very recently [DEF+25]. There, Dolbeault, Esteban, Figalli, Frank and Loss obtained sharp stability results for the Sobolev inequality with optimal dimensional dependence (see also, [CLT+25]). In our case, the inequality (3.2) provides not only stability for the $L^{p}$-Poincaré inequality, but also gives an explicit geometric dependence of the stability constant due to the result of Ferone, Nitsch and Trombetti [FNT12]. Additionally, we show some corresponding results for the Gaussian probability measure using the same logic thanks to [CQS25].
@@ 3. Main results
The identity, as with all such identities, thanks to Lemmata 2.6, 2.7 and 2.8 allows us to derive all possible optimizers of the $L^{p}$-Poincaré inequality for any $p\in\left(1,\infty\right)$. Such set, we will denote as $E_{Poin}:=\{cu_{1}:c\in\mathbb{C}\}$. If we restrict the $L^{p}$-Poincaré inequality to real-valued functions, then the set of all optimizers becomes $E_{Poin}=\{cu_{1}:c\in\mathbb{R}\}$.
@@
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx\geq c_{1}(p)\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}d(u,E_{Poin})^{p}.$ (3.2)
@@
$\pi_{p}=2\int_{0}^{+\infty}\frac{1}{1+\frac{1}{p-1}s^{p}}ds=2\pi\frac{(p-1)^{1/p}}{p(\sin(\pi/p))}\quad,\\ \quad c_{1}(p)=\inf_{(s,t)\in\mathbb{R}^{2}\setminus\{(0,0)\}}\frac{\bigl[t^{2}+s^{2}+2s+1\bigr]^{\frac{p}{2}}-1-ps}{\bigl[t^{2}+s^{2}\bigr]^{\frac{p}{2}}}\in(0,1]$
@@
When $p=2$, in (3.2), interestingly the constant $c_{1}(2)=1$ and $\pi_{2}=\pi$, which implies:
@@
$\lambda_{2}(p,\Omega)-\lambda_{1}(p,\Omega)\geq c_{1}(p)\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}C(p,\Omega,u_{1},u_{2}),$ (3.4)
@@
When $p=2$, in the inequality (3.4), we have that $C(2,\Omega,u_{1},u_{2})=1=c_{1}(2)$ and $\pi_{2}=\pi$, which gives us the fundamental gap of the Laplacian in the form of Yu, Zhong and Smits [YZ86, SMI96].
@@
$\int_{\Omega}|\nabla u|^{p}d\gamma-\lambda_{1}(p,\Omega,\gamma)\int_{\Omega}|u|^{p}d\gamma\geq c_{1}(p)\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}d(u,E_{Poin}^{\gamma})^{p},$ (3.7)
@@
$\pi_{p}=2\int_{0}^{+\infty}\frac{1}{1+\frac{1}{p-1}s^{p}}ds=2\pi\frac{(p-1)^{1/p}}{p(\sin(\pi/p))}\quad,\\ \quad c_{1}(p)=\inf_{(s,t)\in\mathbb{R}^{2}\setminus\{(0,0)\}}\frac{\bigl[t^{2}+s^{2}+2s+1\bigr]^{\frac{p}{2}}-1-ps}{\bigl[t^{2}+s^{2}\bigr]^{\frac{p}{2}}}\in(0,1]$
@@
Finally, applying the bound on the constant $C_{\Omega,p,\omega}$ from Theorem 2.4, we complete the proof. ∎
@@
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx\geq c_{1}(p)\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}\inf_{c\in\mathbb{R}}\int_{\Omega}\left|u-cu_{1}\right|^{p}dx.$ (4.3)
@@
$\lambda_{2}(p,\Omega)-\lambda_{1}(p,\Omega)\geq c_{1}(p)\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}\inf_{c\in\mathbb{R}}\int_{\Omega}\left|u_{2}-cu_{1}\right|^{p}dx,$
@@
5. Proofs for the Gaussian probability measure–Proofs of Theorems 3.10 and 3.12
@@
Applying the lower bound of the constant $C_{\Omega,p,\omega}$ from Theorem 2.4, we complete the proof. ∎
Acknowledgments. We would like to thank Vladimir Bobkov for helpful comments and for pointing out an important reference [FT02] in the field.

### v3

# arXiv:2602.05968v3[math.AP] 28 Feb 2026

@@ Abstract.
In this paper, we obtain stability results for the $L^{p}$-Poincaré inequality for both Lebesgue and Gaussian probability measures (Theorem 3.3 and Theorem 3.13) that involve explicit dependence on the geometry of the domain. As a byproduct, the explicit constant allows us to recover important results of Yu, Zhong [YZ86] and Smits [SMI96] (Corollary 3.9), related to the fundamental gap conjecture of the Laplacian (resolved by Andrews and Clutterbuck [AC11]), thereby providing an alternative proof. Moreover, we extend this spectral gap result to the $p$-Laplacian (Corollary 3.6). Such gap estimates for the Dirichlet $p$-Laplacian appear to be unavailable, as also observed in [DSW18]. Our approach relies on properties of the first eigenfunction of the (Gaussian) $p$-Laplacian operator and weighted Poincaré inequalities for log-concave measures on convex domains.
@@
Poincaré inequality, stability inequalities, Gaussian Poincaré inequality, spectral gap
@@ 1. Introduction
and $\langle\cdot,\cdot\rangle$ is the usual scalar product in $L^{2}$. It is important to note that the result (1.3), in [FT02], was then used to obtain an existence result concerning a (p1)(p-1)-homogeneous problem with Dirichlet $p$-Laplacian.
@@
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}d(u,E_{Poin})^{p}.$ (1.5)
@@
$\pi_{p}=2\int_{0}^{+\infty}\frac{1}{1+\frac{1}{p-1}s^{p}}ds=2\pi\frac{(p-1)^{1/p}}{p(\sin(\pi/p))}$
@@
The constant, in (1.5), is explicit with clear dependence on the diameter of the convex domain $\Omega$, which we believe is the first such result. In general, in most stability results, the explicit form of the constant that appears nearby the distance function (1.1) is unknown. Even for the historic result of Bianchi and Egnell [BE91] for the Sobolev inequality, which was obtained in 1991, no information was known for the constant $c_{S}$ (except an upper bound [KÖN23]) until very recently [DEF+25]. There, Dolbeault, Esteban, Figalli, Frank and Loss obtained sharp stability results for the Sobolev inequality with optimal dimensional dependence (see also, [CLT+25]). In our case, the inequality (3.2) provides not only stability for the $L^{p}$-Poincaré inequality, but also gives an explicit geometric dependence of the stability constant due to the result of Ferone, Nitsch and Trombetti [FNT12]. Additionally, we show some corresponding results for the Gaussian probability measure using the same logic thanks to [CQS25].
Moreover, after normalizing the eigenfunctions $u_{1}$ and $u_{2}$, we obtain the following spectral gap of the Dirichlet $p$-Laplacian:
$\lambda_{2}(p,\Omega)-\lambda_{1}(p,\Omega)\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}C(p,\Omega,u_{1},u_{2}),$ (1.6)
@@
We could not find any results related to the spectral gaps for the Dirichlet $p$-Laplacian operator. In fact, in [DSW18, Page 7], Dai, Seto and Wei, note that “similar gap estimate for the $p$-Laplacian is still unknown”. We believe (1.6) is the first such estimate.
@@
The difference between the first two eigenvalues of the Dirichlet Laplacian on a bounded open domain $\Omega$, denoted by $\lambda_{2}(p,\Omega)-\lambda_{1}(p,\Omega)$, is commonly known as the fundamental gap of $\Omega$. The fundamental gap problem has an interesting history and we give a brief overview of it here.
Originally, van den Berg, in [VAN83], first observed that many convex sets have a spectral gap of $\frac{3\pi^{2}}{\text{diam}(\Omega)^{2}}$. This was also independently observed and then conjectured to be true in any convex domain $\Omega$ by Yau [YAU87], Ashbaugh and Benguria [AB89] for the Schrödinger operator $\Delta+V$. The general statement in terms of the Schrödinger operator has been resolved in the one-dimensional setting by Lavine [LAV94] (see, also, [AB89, HOR03]). In higher dimensions, a breakthrough was made by Singer, Wong, Yau and Yau [SWY+85], where they obtained a lower bound of $\frac{\pi^{2}}{4\text{diam}(\Omega)^{2}}$ by utilizing the gradient estimates in spirit of those from Li [LI79], Li and Yau [LY80] and the log-concavity of the first eigenfunction (established for convex potentials by Brascamp and Lieb [BL76]). Later, Yu, Zhong [YZ86] and [SMI96] improved the estimate to $\frac{\pi^{2}}{\text{diam}(\Omega)^{2}}$. Under certain assumptions on the domain $\Omega$, major step forward in proving the conjecture was made by Bañuelos and Méndez-Hernández [BM00], Davis [DAV01] and Bañuelos with Kröger [BK01]. Finally, Andrews and Clutterbuck [AC11] proved the conjecture in 2011 for any convex domain $\Omega$. A few years later, an alternative proof was given by Ni in [NI13]. Gap estimates on other domains are addressed in [CO97, OSW99, SWW19, HWZ20, DSW21]. In a very recent paper [ABF24], Amato, Bucur and Fragalà strengthened the result of Andrews and Clutterbuck and quantified it in terms of flatness, thereby answering an additional open problem by Yau [YAU93]. Additional works related to the fundamental gap problem of the Neumann Laplacian are present in [PW60, ZHO84]. We also refer to [AND14, DSW18] for comprehensive surveys on the subject.
@@ 2. Preliminaries
The constant, in (2.7), is especially interesting. In [CKL+24b], the authors asked whether it is the most optimal for the given inequality. This question was recently answered by Huang and Tong [HT25] and we state it here.
Lemma 2.7 ([HT25, Theorem 6]).
Let $p\geq 2$. The sharp constant $c_{1}(p)$ in (2.7) is determined as
$c_{1}(p)=(p-1)(1-k_{0})^{p}+pk_{0}(1-k_{0})^{p-1}+k_{0}^{p}>0,$ (2.8)
where $k_{0}=\frac{r_{0}}{1+r_{0}}$ and $r_{0}$ is the solution of the equation
$r^{p-1}-(p-1)r-(p-2)=0.$ (2.9)
Example 2.8.
For $p=3$, the sharp constant is given by $c_{1}(3)=2-\sqrt{2}$.
Remark 2.9.
Note that since $k_{0}=\frac{r_{0}}{1+r_{0}}$, we have $1-k_{0}=\frac{1}{1+r_{0}}$. Substituting this into (2.8) gives us
$c_{1}(p)=(p-1)\left(\frac{1}{1+r_{0}}\right)^{p}+p\frac{r_{0}}{1+r_{0}}\left(\frac{1}{1+r_{0}}\right)^{p-1}+\left(\frac{r_{0}}{1+r_{0}}\right)^{p}.$
This simplifies to
$c_{1}(p)=\frac{p-1+pr_{0}+r^{p}_{0}}{(1+r_{0})^{p}}.$ (2.10)
From equation (2.9), we get that
$r^{p}_{0}=r^{2}_{0}(p-1)+r_{0}(p-2).$ (2.11)
Now we substitute (2.11) to (2.10):
$c_{1}(p)$ $=\frac{p-1+pr_{0}+r^{2}_{0}(p-1)+r_{0}(p-2)}{(1+r_{0})^{p}}$
$=\frac{p-1+(2p-2)r_{0}+r^{2}_{0}(p-1)}{(1+r_{0})^{p}}$
$=(p-1)(r_{0}+1)^{2-p}.$
This is a much simpler expression for $c_{1}(p)$ and it is the one which we will use throughout the text.
Although Lemma 2.7 makes it possible to compute the sharp constant $c_{1}(p)$ via a simple polynomial equation, it is unclear what happens as $p\to\infty$, which is interesting on its own right. That is why, using Lemma 2.7, we provide lower and upper bounds for the constant $c_{1}(p)$ (for more details see Section 4) that exactly reveal its behavior as $p\to\infty$.
Proposition 2.10.
For any $p\geq 2$, we have
$\frac{1}{2^{p-2}}\leq c_{1}(p)\leq\frac{p-1}{2^{p-2}}.$
Moreover, as $p\to\infty$, the constant $c_{1}(p)\to 0$.
@@ 3. Main results
The identity, as with all such identities, thanks to Lemmata 2.6, 2.11 and 2.12 allows us to derive all possible optimizers of the $L^{p}$-Poincaré inequality for any $p\in\left(1,\infty\right)$. Such set, we will denote as $E_{Poin}:=\{cu_{1}:c\in\mathbb{C}\}$. If we restrict the $L^{p}$-Poincaré inequality to real-valued functions, then the set of all optimizers becomes $E_{Poin}=\{cu_{1}:c\in\mathbb{R}\}$.
@@
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}d(u,E_{Poin})^{p}.$ (3.2)
@@
$\pi_{p}=2\int_{0}^{+\infty}\frac{1}{1+\frac{1}{p-1}s^{p}}ds=2\pi\frac{(p-1)^{1/p}}{p(\sin(\pi/p))}$
@@
When $p=2$, in (3.2), interestingly the constant $\pi_{2}=\pi$, which implies:
@@
$\lambda_{2}(p,\Omega)-\lambda_{1}(p,\Omega)\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}C(p,\Omega,u_{1},u_{2}),$ (3.4)
@@
In the aforementioned survey [DSW18, Page 7], Dai, Seto and Wei comment that spectral gaps for the Dirichlet $p$-Laplacian operator are “still unknown” and we believe it is the first such result in this direction.
Remark 3.8.
@@
When $p=2$, in the inequality (3.4), we have that $C(2,\Omega,u_{1},u_{2})=1$ and $\pi_{2}=\pi$, which gives us the fundamental gap of the Laplacian in the form of Yu, Zhong and Smits [YZ86, SMI96].
@@
$\int_{\Omega}|\nabla u|^{p}d\gamma-\lambda_{1}(p,\Omega,\gamma)\int_{\Omega}|u|^{p}d\gamma\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}d(u,E_{Poin}^{\gamma})^{p},$ (3.7)
@@
$\pi_{p}=2\int_{0}^{+\infty}\frac{1}{1+\frac{1}{p-1}s^{p}}ds=2\pi\frac{(p-1)^{1/p}}{p(\sin(\pi/p))}$
@@
4. Proof of Proposition 2.10
Proof of Proposition 2.10.
Let us recall that
$c_{1}(p)=(p-1)(r_{0}+1)^{2-p},$ (4.1)
where $r_{0}$ is the root of the function
$f(r)=r^{p-1}-(p-1)r-(p-2).$
When $p=2$, we have $c_{1}(2)=1$. When $p>2$, we first take the derivative
$f^{\prime}(r)=(p-1)(r^{p-2}-1).$
For any $p>2$, we see that $f^{\prime}(r)=0$ at $r=1$. In addition, for any $p>2$ and $r>1$, one has $f^{\prime}(r)>0$. Also note that f(1)=42pf(1)=4-2p starts at a negative value and becomes positive for large $r$. By continuity of $f(r)$, there exists at least one root $r_{0}>1$. Now since $f(r)$ is strictly increasing on $r\in(1,\infty)$, by the Intermediate Value Theorem, the root is unique and satisfies $r_{0}>1$.
Since $r_{0}>1$, we have $r_{0}+1>2$ and consequently $(r_{0}+1)^{p-2}>2^{p-2}$. Therefore, from (4.1), we obtain that for $p>2$:
$c_{1}(p)<\frac{p-1}{2^{p-2}}.$
Combining $p=2$ and $p>2$ cases, we get that
$c_{1}(p)\leq\frac{p-1}{2^{p-2}}.$
To get a lower bound, we construct a new function $g(t)=t^{q}$, which is convex for any $q\geq 1$ and $t>0$. Since $g(t)$ is convex, it is also midpoint convex:
$g\left(\frac{r_{0}+1}{2}\right)\leq\frac{g(r_{0})+g(1)}{2},$
which is equivalent to
$\left(\frac{r_{0}+1}{2}\right)^{q}\leq\frac{r^{q}_{0}+1}{2}.$
Applying the inequality for q=p1q=p-1 ($p\geq 2$) and multiplying both sides by $2^{p-1}$, gives us
$(r_{0}+1)^{p-1}\leq(r^{p-1}_{0}+1)2^{p-2}.$
Since $r^{p-1}_{0}+1=(r_{0}+1)(p-1)$, we get
$(r_{0}+1)^{p-1}\leq 2^{p-2}(r_{0}+1)(p-1).$
Dividing both sides by $r_{0}+1$, we obtain
$(r_{0}+1)^{p-2}\leq 2^{p-2}(p-1).$
Therefore, from (4.1), for $p\geq 2$, we have
$c_{1}(p)\geq\frac{1}{2^{p-2}}.$
By the Squeeze Theorem, it follows that $c_{1}(p)\to 0$ as $p\to\infty$. ∎
@@
Finally, applying the bounds on the constants $c_{1}(p)$ and $C_{\Omega,p,\omega}$ from Proposition 2.10 and Theorem 2.4, respectively, we complete the proof. ∎
@@
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}\inf_{c\in\mathbb{R}}\int_{\Omega}\left|u-cu_{1}\right|^{p}dx.$ (5.3)
@@
$\lambda_{2}(p,\Omega)-\lambda_{1}(p,\Omega)\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}\inf_{c\in\mathbb{R}}\int_{\Omega}\left|u_{2}-cu_{1}\right|^{p}dx,$
@@
6. Proofs for the Gaussian probability measure–Proofs of Theorems 3.11 and 3.13
@@
Applying the lower bounds on the constants $c_{1}(p)$ and $C_{\Omega,p,\omega}$ from Proposition 2.10 and Theorem 2.4, respectively, we complete the proof. ∎
Acknowledgments. We would like to thank Vladimir Bobkov for valuable discussions and comments as well as for pointing out an important reference [FT02] in the field.